OpenStudy (anonymous):

Write the equation of the line that passes through (−3, 5) and (2, 10) in slope-intercept form

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -3}}\quad ,&{\color{blue}{ 5}})\quad % (c,d) &({\color{red}{ 2}}\quad ,&{\color{blue}{ 10}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ 10}}-{\color{blue}{ 5}}}{{\color{red}{ 2}}-{\color{red}{ (-3)}}}\implies \cfrac{5}{2+3}\implies \cfrac{\cancel{5}}{\cancel{5}}\implies 1 \\ \quad \\ % point-slope intercept y-{\color{blue}{ 5}}={\color{green}{ 1}}(x-{\color{red}{ (-3)}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y-5=x+3

OpenStudy (jdoe0001):

yeap, then solve for "y" to end up with the slope-intercept form

OpenStudy (anonymous):

y=-10

OpenStudy (jdoe0001):

hmm y-5 +5 = x+3 + 5 <--- adding 5 to both sides y = x+8